# User Experience in Dataset Search Platform Interfaces

### v1

@@ 3.4.4. Use of Stimulated Retrospective Think-Aloud
3.5. Implement Questionnaires
3.5.1. Demographic Questionnaires
@@
Utilised at commencement to gain insights into participant backgrounds (Boynton and Greenhalgh, 2004). These questionnaires capture specifics such as professional skills, educational history, and prior familiarity with dataset search utilities. Comprising seven questions, they targets broad categories like ’General’, ’Education’, and ’Occupation’. It solicits information about age group, gender, highest educational attainment, work or study field, frequency of dataset search tool usage, skill level, and previously used tools. Both multiple-choice and open-ended question types are employed.
3.5.2. Post-Task Questionnaires
@@
Issued upon task completion to collect user opinions (Lewis, 2014). They examine usability facets including user satisfaction, perceived utility, and suggested refinements. These surveys encompass four main areas: ’Initial Interaction’, ’Search Operation’, ’Dataset Study’, and ’Filter and Sort Utilisation’. Questions cover a range of topics, from initial impressions and the relevance of search outcomes to the quality of dataset metadata and the efficacy of additional features. A 1-5 rating scale is used for certain queries, while open-ended questions solicit suggestions for enhancements.
3.6. Participant Recruitment
@@
This study recruited 19 participants from a diverse pool, including both academia and industry, via various channels such as LinkedIn, Boss Zhiping, and university email campaigns. This group included 2 UX designers, 2 product managers, 1 special education aide, and 14 students from fields like Human-Computer Interaction and health data science, with ages ranging from 18 to 40. Prior to this study, ethical approval was obtained and informed consent was collected from all participants. All 19 participants tested both tools with the same task.
@@
This study utilized scenarios involving COVID-19 and social sciences in user tasks, primarily to motivate interaction with the platform, rather than to concentrate on the social or health sciences. The goal of this project was to assess overall user experience, focusing on behavior, functionality, and interface design across diverse search websites. These scenarios were instrumental in prompting users to test specific platform features, such as the autocomplete response to ’COVID-19’ in the search bar or the filter effectiveness in the ’social sciences’ category. Consequently, this research was not centered on deriving conclusions specific to social sciences, leading to the recruitment of participants from a wide array of fields.
3.7. Transcription and Coding Scheme
@@
A single researcher manually transcribed 19 videos, organizing the data into 6 primary themes, including ’Initial Interaction’ and ’Search Process’. The coding process combined top-down and bottom-up approaches as outlined by Gu et al. (Gu, 2014). Preconceived notions about web search biases, such as prior beliefs  (White, 2014), informed some of the top-down codes, whereas bottom-up codes emerged and were adapted based on the analysis of participants’ think-aloud sessions, particularly their interactions with interface elements and their engagement with statistics. Codes were designed to be overlapping to capture the multifaceted nature of user experiences. To confirm reliability, the initial coding was reviewed at a later stage by the same researcher, and any deviations were corrected through iterative revisions. Frequency counts were kept for each code to weigh their impact on participant choices.
4. Results
@@
This section is divided into 3 subsections. The first 2 subsections document feedback for Google Dataset Search and data.europe.eu, labeled ’P’ for positive and ’N’ for negative. Each of these subsections includes six sub-subsections corresponding to the user experience elements identified in the user task. All percentage values have been rounded to two decimal places.
@@
While certain insights, such as interface design elements, are specific to each platform, most of the findings regarding mechanisms and functionalities can be applied more broadly to dataset search platforms. Therefore, the last subsection  4.3 presents the development of 10 new design prototypes, based on observations of features common in dataset search platforms across various fields. In the  4.3.3 Dataset Exploration,  4.3.4 Usage of Filters and Sorting,  4.3.5 Dataset Actions, there’s an enhancement in metadata accessibility. For others, the overall current state of dataset search platforms has been improved.
4.1. Google Dataset Search
4.1.1. Initial Interaction
P1: 68% participants noted the search box, search button, help, and feedback features as easily accessible and user-friendly, commenting that the design was ”extremely neat and comfortable.” P2: 79% participants expressed a clear understanding and appreciation of the layout of the search results page. They found the meaning of each filter to be easily comprehensible and praised the filters’ placement at the top of the page, as well as the well-balanced proportions between the preview and detailed information sections. However, N1: 26% participants observed that the language switch button was not easily noticeable, being located at the bottom left corner. One participant said, ”When I first see this Google Dataset Search page, I think of the traditional Google browser search page, where all function buttons are placed at the top of the page and are immediately visible to users.” N2: 16% participants hovered their cursor over the language switch button, indicating confusion. One participant noted, ”Usually such icons are represented by the flag of the country or are explicitly labelled with the language type, which makes it easy to understand at first glance.” N3: In a multitasking scenario involving a note-taking app, 21% participants found that the detailed dataset information appeared compressed and elongated, requiring constant scrolling. They commented, ”The detailed introduction of the dataset is compressed to the point where only two or three words are displayed in one line. This is unreadable.” N4: 16% participants hovered their mouse over the bookmark icon due to its unclear design. One of them stated, ”It looks like a report button.” N5: 11% participants were unsure about making immediate actions like bookmarking, sharing, or citing a dataset. One said, ”I was looking for a way to mark this dataset, but I couldn’t find it,” after extensive scrolling. N6: An undergraduate participant from mainland China pointed out that switching the website’s language only updated the structure, not the content of dataset details. Both the participant and the researchers acknowledged the limitation of the website’s language features for users of browsers lacking built-in translation.
4.1.2. Search Process
P7: 63% participants start to input search queries sequentially, utilising system suggestions before finalising keywords. This strategy aids in objective-specific result retrieval and avoids ineffective keywords. In some cases, system-provided datasets already met research task requirements. A participant noted, ”This feature sometimes yields unexpected results!” P8: Google Dataset Search recognised misspelled words. For instance, typing ”deth” retrieved datasets related to ”death.” P9: For misspelled terms like ”covai19,” Google Dataset Search failed to retrieve datasets but offered suggestions like ”Make sure that all words are spelled correctly.” P10: 21% Participants effectively utilised Boolean and Exact Phrase operators. Use of quotes made a discernable difference in search results. However, N14: Using ”covid-19 cases” as a prefix didn’t yield UK-specific datasets. Switching to ”covid-19 infections” did, but also increased time spent on it. N15: The researchers found that participants began by entering ”England covid-19 cases”, resulting in two system suggestions, with the second one fitting research task requirements. However, not all users were as fortunate. When ”covid-19 cases England” was entered, Google Dataset Search suggested only one dataset, which wasn’t the best match. The rearrangement of search terms had a substantial effect on the results and the time spent searching. N16: 37% participants spent a significant amount of time downloading a dataset to look at its colmun and correlation values, and then gave up on using the dataset because of the large number of missing values in the file. One participant said, ”Luckily the zip file is not too big. N17: Google Dataset Search misunderstood ”covidd-19 infraction” to mean ”covid-19 infection,” revealing its limitations in contextual understanding. The participant commented, ”Google Dataset Search can recognise my spelling errors well, but sometimes it misunderstands what I am trying to convey.” N18: Different spellings of ”covid-19” yielded varying results. The spelling ”covid-19” was more task-related compared to other variants. N19: Misspellings in multiple words made it challenging for users to correct their queries, especially when they didn’t realise a spelling error. N20: Sentence-like queries yielded fewer results than more structured queries. Both aimed to retrieve similar information. N21: The Exclude Words operator ”-” didn’t function as expected, leading to undesired results. N22: Searching for operator-related help was futile. 21% participant remarked, ”The User Guide page of the website is 404, and now it doesn’t provide me with the corresponding help, which is very confusing and wasting my time.”
4.1.3. Dataset Exploration
P13-P16 Participants first pay attention to the title, then the update time and dataset type. One remarked, ”I trust dataset content over the publisher”. Furthermore, dataset ranking influences choices, especially the top three listings. P17 Participants paid particular attention to the descriptions, citation counts, and providers on the right-hand dataset display section. P18 74% participants were indifferent to the specifics of the licence, but 68% preferred its existence. However, N29 53% participants mentioned that they initially cared a lot about the ranking on the left side of the page. 26% participants said they didn’t care about the ranking at all and often checked many datasets. One participant even scrolled to the bottom of the page, causing the page to refresh twice before they stopped checking datasets. One participant stated, ”I feel that both datasets ranked high and low can meet my search requirements”. N30 21% participants were frustrated by incomplete dataset titles and limited preview information. N31 and N32 26% participants showed a lack of interest in the dataset provider and dense description, commenting ”I only care if the dataset has a provider, not the specific provider name” and ”Descriptions too dense to be read only by machines, not people!” N33 37% participants point out a lack of detailed file information, and N34 16% participants indicated a high implementation difficulty in packing more information into the left preview module. One participant said, ”Although I hope to fit more information, I would still prefer to see a clean preview section on a page.” N35 21% participants after exploring a qualifying dataset, would click on the website link given in ’explore at’ to view more information about the dataset. However, 11% participants stated, ”Currently, I have to click on ’Explore at’ to jump to the corresponding page to see more detailed information, which is somewhat troublesome. I need to switch between several websites, and sometimes there are too many tabs, and I forget which page was transferred from which link.”
4.1.4. Usage of Filters and Sorting
P23 68% participants intuitively understood the names and effects of the filters without requiring additional help. N42 32% participants underlined the difficulty in sorting the datasets by their recency; a participant stated, ”The dataset search results are sorted too simply. I just want to find a recent dataset, but there is no way to do it.” N43 26% participants noticed that for multiple choice filters such as ”topic”, when they chose two options, the ”+1” displayed on the filter could be misleading. 16% of them interpret this ”+1” as indicating that only one ”topic” has been selected. N44 16% participants preferred interactive animations for dropdown symbols for a more intuitive interface experience, articulated by a participant’s statement, ”Since this is a dropdown symbol, why doesn’t it have the animation effect of a dropdown symbol? I thought the webpage was not loaded.” N45 16% participants highlighted concerns about the broad time spans in the ’Last updated’ filter; as one participant noted, ”There are few datasets in the last month, and there are too many datasets in the last year. Isn’t there an intermediate time period to choose?” N46 47% participants exposed the inconvenience of filter resets upon search query modification; one participant stated, ”This makes it uncomfortable to use, I have to re-select the filter every time.” N47 21% participants observed that some datasets lacked usage permission tags, causing issues with filter-based searches; a participant asked, ”Now whether I click on ’commercial use allowed’ or ’Non-commercial use allowed’, only two datasets are always displayed, but in fact, there are eight available datasets in the current search results, so what usage permission category do the other six datasets belong to?” N48 11% participants criticised the broad categorisation within the ’Topics’ filter, highlighted by a participant who said, ”Especially the ’Natural Science’ and ’Engineering’ tags, their scope can basically include the scope of other search tags.” N49 47% participants expressed confusion about the classification of the datasets retrieved after applying the filters. One participant stated, ”I don’t even know which filters this dataset satisfies. I need to read the entire introduction to get a feel for it. In this case, the filters don’t seem to have worked and haven’t helped me save time.” Lastly, N50 11% participants ignored filters altogether; one stated, ”Honestly, I ignored the existence of the filters at the beginning, mainly because the first search page of Google Dataset Search doesn’t have filter options, just a search box.”
4.1.5. Dataset Actions
P25 All participants stated that they easily found and used the ”Save”, ”Share”, ”Cite” buttons when they found a suitable dataset. They also reported that it was very simple to locate previously saved datasets through the ”Saved datasets” button. Many participants expressed, ”The position of these three buttons is very clear, and I could see them at first glance”. Yet, N56 16% participants preferred to use their browser’s built-in share function. One participant said, “Compared to directly copying the link from the top of the browser, clicking the share button and then choosing the sharing method seems very complicated.” N57 The researcher asked all participants ”Would you use these four sharing methods? Why?” 37% participants stated they would not use the Facebook and Twitter sharing options. One user stated, “I generally only use the dataset search website in a work context”. Moreover, 21% participants stated, ”I would share this dataset via email, but typically, the dataset is only a small part of our email content. Therefore, I wouldn’t write an email solely because of this dataset; I would reference this dataset where needed in the email, so I wouldn’t use the current email sharing feature.”. N58 21% participants misunderstood the ”Click to copy link” as merely informational text rather than an interactive button, attempting to select the URL beneath it instead. N59 16% participants found that the website does not specify the current citation format and standard. One participant mentioned, “I basically would not directly use the format provided by this citation, because it doesn’t even tell me what format it uses.”
4.1.6. Feedback and Help
P27: 32% participants pointed out that google provides what is a dataset search and how to provide a dataset, which is very friendly to people without basic knowledge. P28: 84% participants easily found the ”Help” and ”Feedback” buttons, and two participants independently browsed the content in the ”Help” page at the beginning. N62: 11% participants found that the information on the user guide page currently shows a ’404 Page not found’ error, and participants need to jump to multiple pages in the help documentation to understand Google Dataset Search, which adds a lot of reading time.
4.2. data.europa.eu
4.2.1. Initial Interaction
P3: 11% participants commended the comprehensive summary of datasets on the initial page of data.europa.eu. One participant cited the inclusion of datasets from 36 countries as instilling confidence, while another was impressed by the sheer number of datasets, stating, ”over 1,000,000 datasets shocked me a lot, I feel this website must contain a suitable dataset for my requirement.” P4: The display of a page path was seen as useful by 26% participants for quick navigation. P5: A window hover and shadow animation were appreciated by 42% participants. One stated, ”compared to Google Dataset Search, this animation makes me feel more clear and conspicuous about what I have chosen.” P6: The visibility of selected filters at the top of the results preview section was praised by 26% participants for providing an overview. One remarked, ”Especially with multi-select filters, I can instantly see all the options I’ve selected, which indeed makes subsequent modifications more convenient for me.” However, N7: 32% Participants using smaller laptop screens criticized the absence of a readily visible search box. One participant stated, ”If I didn’t know at first that this web is a dataset search tool, I wouldn’t believe it has a search bar, because I can’t see it at first glance when I enter this website, and this search engine design style is not common.” N8: 53% participants reported distractions due to unrelated windows. One mentioned, ”I came to search for datasets, not to see any advertisements or events”, and another criticized the switching advertisements, stating, ”The switching always disturbs my thought process when I am thinking about what keywords to put into this search bar.” N9: A UX designer noted the absence of indicators for the search bar, making it appear non-functional. N10: 37% participants reported cumbersome scrolling to select filters, which also shifted positions upon selection. One participant expressed frustration, stating, ”The process of selecting filters is really uncomfortable. I need to scroll the page back and forth many times to complete the selection, and why does the position of the selected filter change?” N11: 11% participants found it inconvenient that text in the preview section couldn’t be copied for note-taking. One said, ”I have no choice but to open this dataset, copy the information I need from its detailed introduction page, which is a waste of time.” N12: Although colour differentiation for tags was appreciated, it was found to be inaccessible for individuals with different types of colour blindness. N13: 11% participants criticized the ’show more’ button as redundant when there’s minimal hidden content. They commented, ”In order to see more content, I need to click the ’show more’ button, but the hidden content is minimal, so I feel this click interaction is meaningless.”
4.2.2. Search Process
P11: 63% participants found the ability to sort datasets by ”Last modified” and ”Last created” beneficial when they had time constraints. One participant stated, ”The feature to sort by release time is very useful, especially after all keywords and filters have been selected, I just want to see the latest published datasets under all the conditions I have set.” P12: 21% participants found the separate search options for catalogues and datasets provided in the search bar to be very effective. One participant stated, ”Catalogues provide me with information about publishers from a specific country, and I can find the dataset I need by searching for keywords in all the datasets published by this publisher, which greatly improves my search efficiency.” N23: 26% participants observed that, unlike Google dataset search, more strings entered in the search box resulted in a broader search scope instead of narrowing it down. Many participants stated, ”I would prefer the search system to narrow down the search scope based on my search strings, rather than continue to expand it.” N24: 16% participants noted that sorting by a single criterion led to extreme results, especially when many datasets remained. One participant stated, ”I hope the priority in dataset searches is often to find the most relevant first and then find the most recent datasets.” N25: 16% participants noticed the absence of transition animations for the dropdown symbols on the right side of ”Datasets” and ”Sort by”. After clicking ”More” or ”Less filter”, the dropdown symbol changes significantly, indicating the unfolding and folding of results. One participant stated, ”This inconsistency in the overall design can mislead me a lot.”
4.2.3. Dataset Exploration
P20 63% participants indicates strong appreciation for the use of colour tags to denote dataset types, as one participant said, ”This design makes me feel very comfortable, and the types of files are clear at a glance.” P21 42% participants found the similar dataset feature useful when the similarity was high, capturing the sentiment with a comment, ”I didn’t expect that this very similar dataset was published even more recently than the dataset I found.” P22 21% participants discusse the utility of selecting ’Items per Page’, elaborating with a participant’s comment, ”I think displaying ten datasets at a time is very few, and I always have to click next page to refresh new datasets, which I don’t want to wait for.” N36 47% participants suggest the possible improvement of using colour tags for keyword matching, as multiple participants said, ”I would prefer these different colour tags to be about keyword information.” N37 37% participants signify the importance of ’Update Time’ and other dataset metadata like size and missing values ratio, summarised by a participant’s comment, ”as these factors influence whether the dataset warrants further examination.” N38 21% participants highlight the emphasis users put on ’Publisher’ and ’Updated’ in the expansion box, but found ’Contact,’ ’Identifiers,’ etc., less significant. One participant noted, ”I don’t need to understand the contact of a dataset at all, my primary purpose is to check whether the metadata of the dataset is suitable.” N39 42% mentions the confusion over non-English keywords and the difficulty in identifying query-related keywords from the list, summarised by one comment, ”This dataset covers too many keywords, and I just want to see at a glance which keywords from my search strings are included.” N40 16% participants expose user difficulty in understanding the ’Less Similar’ tag, with one participant stating, ”Instead of recording similarity here, why not directly state which keywords are included.” N41 16% participants found that datasets with the ”less similarity” tag were generally far from their target dataset. They found that only datasets with the ”similar” tag might be worth further exploration.
4.2.4. Usage of Filters and Sorting
P24 74% participants stated that the dual provision of filtering and sorting methods on the website was lauded for enhancing efficiency in dataset discovery. However, N51 58% participants exposed a time-consuming issue, where participants found that the page refreshes after each filter selection, especially problematic when applying multiple filters. N52 37% participants thought that refreshing the entire page after each filter selection could mislead users into thinking that the filter options are for single selection only, even though some filter options allow multiple selections, potentially resulting in imprecise search results. One participant, after selecting an option and noticing the page refresh and all expanded options collapse, skipped this filter and stated, ”It would be better if it was multi-select,” even though the filter was indeed multi-select. N53 16% participants spotlighted another issue where three participants were unaware that a filter was for single selection until they noticed that their previous selection was overwritten by the new one; a participant stated, ”I thought it was multi-select, and I wouldn’t have known it was single-select if the website’s upper right side didn’t display all my selected filter options.” N54 11% participants noted that the presence of numbers in the text of some filter options, such as ”Version 2”, can occasionally impact their interpretation of the number of datasets. One participant stated, ”At first glance, I thought there were only two datasets that met this filter requirement,” due to the confusion between the dataset count and the numerical information within the filter option’s text. Lastly, N55 26% participants found it easy to overlook the search box within each filter. One participant stated, ”as the placeholder text in the search box says ’Filter’, leading me to assume that it’s a description of the drop-down menu.”
4.2.5. Dataset Actions
P26 42% participants stated that the citation function was clearly visible and comprehensive, offering four different citation styles, which saved them significant time in creating citations. One participant stated, ”This has saved me a lot of time in creating citations.” N60 11% participants highlighted the absence of a direct LaTeX citation option, stating that manual LaTeX citation creation still constitutes a burdensome process. N61 58% participants stated that when they want to compare similarities between datasets, they need to manually copy and paste dataset information into a separate document, which they find cumbersome. One participant stated, ”It would be great if datasets could be compared directly on the website.”
4.2.6. Feedback and Help

### v2

@@ 3.4.4 Use of Stimulated Retrospective Think-Aloud
3.5 Implement Questionnaires
3.5.1 Demographic Questionnaires
@@
Utilised at commencement to gain insights into participant backgrounds [4]. These questionnaires capture specifics such as professional skills, educational history, and prior familiarity with dataset search utilities. Comprising seven questions, they targets broad categories like ’General’, ’Education’, and ’Occupation’. It solicits information about age group, gender, highest educational attainment, work or study field, frequency of dataset search tool usage, skill level, and previously used tools. Both multiple-choice and open-ended question types are employed.
3.5.2 Post-Task Questionnaires
@@
Issued upon task completion to collect user opinions [18]. They examine usability facets including user satisfaction, perceived utility, and suggested refinements. These surveys encompass four main areas: ’Initial Interaction’, ’Search Operation’, ’Dataset Study’, and ’Filter and Sort Utilisation’. Questions cover a range of topics, from initial impressions and the relevance of search outcomes to the quality of dataset metadata and the efficacy of additional features. A 1-5 rating scale is used for certain queries, while open-ended questions solicit suggestions for enhancements.
3.6 Participant Recruitment
@@
This study involved 19 participants from diverse backgrounds, including academia and industry, recruited through LinkedIn, Boss Zhiping, and university email campaigns. This group consisted of 2 UX designers, 2 product managers, 1 special education aide, and 14 students specializing in Human-Computer Interaction and health data science, with ages ranging from 18 to 40. Ethical approval was obtained, and informed consent was collected from all participants. All participants tested both tools using the same task.
@@
This study used scenarios related to COVID-19 and social sciences to motivate user interaction with the platforms, focusing on overall user experience, behavior, functionality, and interface design across various search websites. The scenarios prompted users to test specific platform features, such as autocomplete responses to ’COVID-19’ and filter effectiveness in the ’social sciences’ category. The research did not aim to draw conclusions specific to social sciences, which is why participants were recruited from a wide range of fields.
3.7 Transcription and Coding Scheme
@@
A single researcher manually transcribed 19 videos, organizing the data into 6 primary themes, including ’Initial Interaction’ and ’Search Process’. The coding process combined top-down and bottom-up approaches as outlined by Gu et al. [13]. Preconceived notions about web search biases, such as prior beliefs  [25], informed some of the top-down codes, whereas bottom-up codes emerged and were adapted based on the analysis of participants’ think-aloud sessions, particularly their interactions with interface elements and their engagement with statistics. Codes were designed to be overlapping to capture the multifaceted nature of user experiences. To confirm reliability, the initial coding was reviewed at a later stage by the same researcher, and any deviations were corrected through iterative revisions. Frequency counts were kept for each code to weigh their impact on participant choices.
4 Results
@@
This section is divided into 3 subsections. The first 2 subsections document feedback for Google Dataset Search and data.europe.eu, labeled ’P’ for positive and ’N’ for negative. Each of these subsections includes six user experience elements identified in the user task. All percentage values have been rounded to two decimal places.
@@
While certain insights, such as interface design elements, are specific to each platform, most of the findings regarding mechanisms and functionalities can be applied more broadly to dataset search platforms. Therefore, the last subsection  4.3 presents the development of 10 new design prototypes, based on observations of features common in dataset search platforms across various fields.
4.1 Google Dataset Search
In the initial interaction, 68% of participants found the search box, button, help, and feedback features accessible and user-friendly, describing the design as "neat and comfortable" (P1). 79% appreciated the layout of the search results page, understanding the filters and praising their placement and the balance between preview and detailed information sections (P2). However, 26% noted the language switch button was hard to find at the bottom left, and 16% found its design confusing, suggesting the use of country flags or explicit labels (N1, N2). In multitasking scenarios, 21% found the detailed dataset information compressed and unreadable (N3).
During the search process, 63% of participants used system suggestions to refine keywords, improving result relevance (P7). Google Dataset Search effectively corrected misspellings but sometimes failed to retrieve datasets for certain terms, offering spelling checks instead (P8, P9). 21% effectively used Boolean and Exact Phrase operators, although specific prefixes did not yield desired results (P10, N14, N15). 37% abandoned datasets due to missing values (N16), and the system sometimes misinterpreted queries (N17). Different spellings and sentence-like queries affected result relevance, and the Exclude Words operator did not function as expected (N18-N21). Searching for help with operators was unproductive as the User Guide page was missing, causing confusion (N22).
In exploring datasets, participants focused on the title, update time, and dataset type, with a preference for content over the publisher (P13-P16). Ranking, especially in the top three listings, influenced choices. Participants valued descriptions, citation counts, and providers in the right-hand display section (P17). While 74% were indifferent to license specifics, 68% preferred its existence (P18). However, 53% initially cared about ranking, though 26% disregarded it and checked many datasets, sometimes scrolling to the page’s bottom (N29). Incomplete titles, limited preview information, and dense descriptions were common frustrations (N30-N32). 37% highlighted a lack of detailed file information (N33), and despite difficulties in packing more information into previews, 16% still preferred a clean preview section (N34). After finding a suitable dataset, 21% clicked on ’explore at’ for more details, but 11% found this process cumbersome due to switching between multiple tabs (N35).
Regarding filters and sorting, 68% intuitively understood filter names and effects (P23). However, 32% found it difficult to sort datasets by recency (N42). Multiple-choice filters like "topic" confused 26% of participants due to the misleading "+1" display, with 16% misinterpreting it as indicating only one topic selected (N43). 16% preferred interactive animations for dropdown symbols, and 16% found the broad time spans in the ’Last updated’ filter problematic (N44, N45). 47% were inconvenienced by filter resets upon search query modification (N46). Issues with datasets lacking usage permission tags and broad categorization within the ’Topics’ filter were noted (N47, N48). 47% expressed confusion about dataset classification after applying filters (N49), and 11% ignored filters altogether due to their initial invisibility (N50).
In terms of dataset actions, all participants easily found and used the "Save," "Share," and "Cite" buttons (P25). However, 16% preferred using their browser’s built-in share function (N56). When asked about sharing methods, 37% wouldn’t use Facebook and Twitter options, preferring work-related contexts (N57). Additionally, 21% found the dedicated email sharing feature less useful (N57). 21% misunderstood the "Click to copy link" as informational text rather than an interactive button (N58), and 16% noted that the website does not specify the current citation format and standard (N59).
Lastly, in feedback and help, 32% appreciated explanations on dataset search and provision (P27). 84% easily found the "Help" and "Feedback" buttons, with some browsing the "Help" page content independently (P28). However, 11% encountered a ’404 Page not found’ error on the user guide page, increasing their reading time due to navigating multiple help pages (N62).
4.2 data.europa.eu
In the initial interaction with data.europa.eu, participants had mixed experiences. 11% appreciated the comprehensive dataset summary and inclusion of datasets from 36 countries (P3). The page path display and window hover animations were positively noted by 26% and 42% of participants, respectively (P4, P5). However, 32% of users on smaller laptop screens found the search box hard to spot (N7), and 53% were distracted by unrelated windows and advertisements (N8). A UX designer highlighted the absence of search bar indicators (N9), and 37% found filter selection cumbersome due to shifting positions (N10). Some participants (11%) were frustrated by the inability to copy text from the preview section (N11), and color differentiation for tags was problematic for those with color blindness (N12). The ’show more’ button was seen as redundant by 11% of users (N13).
During the search process, 63% found sorting by "Last modified" and "Last created" useful, especially under time constraints (P11). Separate search options for catalogues and datasets improved search efficiency for 21% of participants (P12). However, 26% noted that more search strings broadened the search scope instead of narrowing it (N23). Sorting by a single criterion led to extreme results for 16% (N24), and the absence of transition animations for dropdown symbols was noted by the same percentage of users (N25).
In dataset exploration, 63% of participants appreciated the use of color tags for dataset types (P20), and 42% found the similar dataset feature beneficial (P21). The option to select ’Items per Page’ was useful for 21% (P22), although 47% suggested using color tags for keyword matching instead (N36). Metadata like ’Update Time’ and size were emphasized by 37% as important for dataset assessment (N37). Confusion over non-English keywords and the ’Less Similar’ tag was mentioned by 42% and 16% of participants, respectively (N39, N40). Some users found that datasets with the "less similarity" tag were not relevant (N41).
Regarding filters and sorting, 74% praised the dual provision of filtering and sorting methods (P24). However, 58% found page refreshes after each filter selection time-consuming (N51). This led to confusion for 37%, as users thought filters were single-selection only (N52). The need for clearer indication of single-selection filters was noted by 16% of participants (N53), while 11% were confused by numbers in filter options’ text (N54). Additionally, 26% found the search box within each filter easy to overlook (N55).
For dataset actions, 42% of participants found the citation function visible and comprehensive, offering multiple citation styles (P26). However, 11% missed a direct LaTeX citation option (N60), and 58% found comparing datasets cumbersome without a direct comparison feature on the website (N61).